OpenStudy (anonymous):

Solve for n 4 = log(2^n -1)

OpenStudy (anonymous):

\[4 = \log(2^n-1)\]

OpenStudy (anonymous):

change it to exponential form than it is a lot easier to solve

OpenStudy (anonymous):

explain please?

OpenStudy (anonymous):

ok logarithmic forms and exponential are interchange able i will show you

OpenStudy (anonymous):

\[y=b^x, x=\log_b(y)\]

OpenStudy (anonymous):

with the way you have your equation your b is equal to 10

OpenStudy (anonymous):

so your equation will be...?

OpenStudy (anonymous):

so I could say, n = -1 (4)

OpenStudy (anonymous):

\[2^n-1=4^{10}\]

OpenStudy (anonymous):

try to solve what I gave you

OpenStudy (anonymous):

I forget logs, I'm not sure how to get the n out.

OpenStudy (anonymous):

ok the formula can be reduced to \[2^n=1048575\]

OpenStudy (anonymous):

log of 1048575 = 2 ?

OpenStudy (anonymous):

oops sorry i mad one mistake is 104857

OpenStudy (anonymous):

um not quite

OpenStudy (anonymous):

log_2(n)=104857

OpenStudy (anonymous):

\[\log_2(n)=104857\] \[\frac{ \log(n) }{ \log(2) }=104857\]

OpenStudy (anonymous):

multiply both sides by log(2)

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry, I'm trying to understand :-\

OpenStudy (anonymous):

no problem it takes time

OpenStudy (anonymous):

when there is a base of a number that is not ten you can split them like how i did

OpenStudy (anonymous):

could you walk me through my problem? \[80 = 20 \log (2^n - 1)\]

OpenStudy (anonymous):

this only works for logs not lns and sure

OpenStudy (anonymous):

divide both sides by 20

OpenStudy (anonymous):

so first I divide by 20 and got what I said earlier.. \[4= \log(2^n - 1)\]

OpenStudy (anonymous):

good now change to exponential i put up the formula. It took me a while to memorize that too so dont worry

OpenStudy (anonymous):

now what is the base? 10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when a base is not written you automatically assume that it is 10

OpenStudy (anonymous):

I can't simplify it more first?

OpenStudy (anonymous):

2^n-1=4^10 should be the formula

OpenStudy (anonymous):

no you have to change it to exponential

OpenStudy (anonymous):

you can't simplify more because everything on the other side is in a log

OpenStudy (anonymous):

so now I have to use log again, but this time base 2?

OpenStudy (anonymous):

yes but when that occurs you can use this formula where x is any number but 10 \[\log_b(y)=\frac{ \log(y) }{ \log(b) }\]

OpenStudy (anonymous):

how are you so far

OpenStudy (anonymous):

you should have gotten \[\log(n)=\log(2)(4^{10}-1)\]

OpenStudy (anonymous):

oops plus 1

OpenStudy (anonymous):

\[\log(n)=315653.129763\]

OpenStudy (anonymous):

now change back to exponential and you are done

OpenStudy (anonymous):

um r you good?

OpenStudy (anonymous):

sorry, minor emergency.

OpenStudy (anonymous):

oh ok no problem

OpenStudy (anonymous):

n = \[n= \frac{ \log(4+1) }{\log 2 }\]

OpenStudy (anonymous):

I was doing something like that

OpenStudy (anonymous):

but I think thats wrong too :-\

OpenStudy (anonymous):

n = 10^315652.5277 ?

OpenStudy (anonymous):

that can't be right...